# Temporal Variation of the Coronal Parameter in a Jetted Tidal Disruption Event: Swift J1644+57

### 10 Mar 2026 v1

@@ IV.1 Nuclear Properties
The Optxagnf model provides several parameters of the source. Among them, we find that the luminosity, in terms of $\log(L/L_{Ed})$, is highly sensitive to the fitting. The fitted spectra exhibited a variation of the order of 2 with the range of our observations. During the initial days, Swift J1644+57 had a bolometric luminosity of around 500 times the Eddington limit. In the entire observation, the source luminosity was in the super-Eddington regime. The luminosity followed a steady decrease from the initial days, with occasional variation in between, which can be associated with the intrinsic photon count variation in the light curve. During XRT17, the last observation, the source luminosity decreased to nearly 3 times the Eddington limit. We find the source luminosity weakly correlated with the hydrogen column density, with a Pearson Correlation coefficient of -0.31.
The spin of Swift J1644+57, measured using the aa* parameter of Optxagnf model, exhibited a region of an intermediate spinning black hole. We find aa* to vary from 0.4 to 0.8. The value of the spin aligns with the earlier claim by Tchekhovskoy et al. (2014), where they predicted the spin of the source should be greater than $\geq 0.5$.
@@ IV.2 Coronal Variations
To measure the coronal size variation of Swift J1644+57, we employ the Optxagnf model to determine the spectral fits using the stacked Swift /XRT data. Through our spectral fitting, we find that at the peak luminosity of Swift J1644+57, the radial size of the corona was $31\pm 2.3~r_{\rm g}$. Subsequently, an almost similar size of the coronal parameter ($R_{cor}$) was found in the XRT2 and XRT3 observational epochs. The size of the corona increases $35\pm 3.6~r_{\rm g}$ in the XRT4 epoch. Later on, in the XRT5 and XRT6 epochs, the $R_{cor}$ keeps increasing to $43\pm 3.7~r_{\rm g}$ and $51\pm 4.3~r_{\rm g}$, respectively. During the XRT7 observation, the $R_{cor}$ is $58\pm 4.1~r_{\rm g}$. Up to the XRT7 observation, we observed a sharp rise in the $R_{cor}$. The fitted value of the $R_{cor}$ is $61\pm 4.2~r_{\rm g}$ at XRT8 and drops to $58\pm 4.5~r_{\rm g}$ at XRT10. From the XRT10 to XRT17 observations, the $R_{cor}$ varies little from $58\,r_{\rm g}$ to $64\,r_{\rm g}$, indicating a saturation level of the parameter. All estimated values of the coronal parameters are depicted in Fig. 8. In addition, we find through our analysis that the $R_{cor}$ is anti-correlated with the luminosity $\log(L/L_{\rm Edd})$ with the Pearson correlation coefficient: 0.87-0.87.
@@
$l_{I}$, $l_{C}$, and $l(t)$ are the initial, final, and the time-dependent Eddington ratios, respectively. Starting with an initial coronal size $R_{I}$, the corona expands following the decay pattern of the luminosity profile and finally saturates to a value of $R_{C}$. We adopt $R_{I}=30\,r_{\rm g}$ and $R_{C}=65\,r_{\rm g}$. The normalized disk luminosity, $\log(L/L_{\rm Edd})$, at each observational epoch is displayed in Table 4. The predicted coronal size at each epoch using Equation (8) is denoted with red circles in Fig. 8. The error in the theoretical values is a contribution arising from the observational error in the measurement of luminosity as presented in Table 4. We find that the evolution of theoretically speculated corona size agrees with that of the observational $R_{cor}$ parameter in Swift J1644+57, and considering the error bars, their variation remains similar.
Figure 6: Optxagnf fitted spectral parameters variation with each other of Swift J1644+57 are presented. Panel (a) shows the variation of $N_{H}$ with respect to the normalized luminosity. The variation of $N_{H}$ with respect to $\Gamma$ is shown in panel (b). Panel (c) demonstrates the weak correlation between $\Gamma$ and scattering fraction $f_{pl}$. Variation of the $\Gamma$ and $\tau$ can be observed in panel (d). Panel (e) presents the variation of coronal radius and spin parameter. Uncorrelated variation between spin and coronal radius can be observed in panel (e).
@@
Figure 7: Markov Chain Monte-Carlo plot for parameter variations of XRT2 and XRT15 are presented. Goodman-Weare (EMCEE) (Foreman-Mackey et al., 2013) is opted for 32 walkers and a chain length of 1 million iterations, where the transient phase is removed by burning the chain up to 9984 iterations. The spin parameter was kept frozen at $a=0.60$ to perform the MCMCs.
@@ IV.2.1 Viability of TDE Corona Models
It should be noted that, to date, there exists no publicly available model in XSPEC to fit the hyper accretion regime of Swift J1644+57. The source was one of the very few jetted TDEs ever observed with a QPO feature. The initial accretion rate was $\sim 1000~L_{Edd}$ and was widely explored for its unique attributes. While choosing the model, one must be extremely cautious. Optxagnf is the only model in XSPEC that applies to the super-Eddington accretion regime of the AGNs, and measures the coronal size.
@@
Recently, Chakraborty et al. (2026) has proposed a coronal emission for TDEs (see their Fig. 7) where the optical depth could be higher, accompanied by a lower electron temperature. The parameters obtained from the Optxagnf model, presented in Table 4, exhibit a similar nature of corona. Although the morphology of the TDE corona could depart from the structure of the warm corona, the thermal properties, such as high $\tau$ and $kT_{s}$, resemble the warm corona of the Optxagnf model. The variation of coronal radius could be observed from the MCMC plots presented in the Fig. 7. However, it should be noted that Optxagnf , developed for AGNs, may not be physically self-consistent for hyper-Eddington jetted TDEs, such as for Swift J1644+57.
@@
Table 4: Optxagnf fitted spectral parameter are presented. We used stacked Swift/XRT over XMM-Newton as the former provides more degrees of freedom over the latter.
ID MJD $N_{H}$ $\log(\frac{L_{o}}{L_{E}})$ $a^{*}$ $r_{cor}$ $kT_{s}$ $\tau$ $\Gamma$ $f_{pl}$ $dof/chi^{2}$
@@
Figure 9: Edge-on view schematic diagram of coronal size variation in the early and late stages of Swift J1644+57. The system comprises a disk around the SMBH, a corona on the disk, and a jet. The disk size is twice the tidal disruption radii, which is approximately estimated to be $2R_{\rm t}\sim 100\,r_{\rm g}$ for the $10^{6}M_{\odot}$ SMBH and a solar-type star (see equation 1).
@@ V Summary
    By interpreting the hard X-ray-emitted region as a corona, we have estimated the size of the corona as a function of time. We confirm that the coronal parameter increases over time for the observed period. We also find that the evolution of the corona size agrees with the theoretical conjecture made by Mummery and Balbus (2021).
@@
The authors acknowledge an anonymous reviewer for their constructive comments. AC acknowledges the UPES Seed grant (UPES/R&D/SoAE/25062025/23) for partially supporting this research. The research of A.C., S.H., and S.S.H. were supported by the Canadian Space Agency (CSA) and the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grants and the Canada Research Chairs programs. The Basic Science Research Program has supported K.H. through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (2016R1A5A1013277 (K.H. and A.C.), 2020R1A2C1007219, and RS-2025-23323627 (K.H.)). K.H. thanks Andrew Mummery for fruitful discussions. AJ acknowledges support from the Fondecyt fellowship (Proyecto 3230303). Research work at the Physical Research Laboratory, Ahmedabad, is funded by the Department of Space, Government of India. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France.

### 07 Apr 2026 v2

@@ IV.1 Nuclear Properties
The Optxagnf model provides several parameters of the source. Among them, we find that the luminosity, in terms of $\log(L/L_{Edd})$, is highly sensitive to the fitting. The fitted spectra exhibited a variation of the order of 2 with the range of our observations. During the initial days, Swift J1644+57 had a bolometric luminosity of around 500 times the Eddington limit. In the entire observation, the source luminosity was in the super-Eddington regime. The luminosity followed a steady decrease from the initial days, with occasional variation in between, which can be associated with the intrinsic photon count variation in the light curve. During XRT17, the last observation, the source luminosity decreased to nearly 3 times the Eddington limit. We find the source luminosity weakly correlated with the hydrogen column density, with a Pearson Correlation coefficient of -0.31.
The spin of Swift J1644+57, measured using the $a^{*}$ parameter of Optxagnf model, exhibited a region of an intermediate spinning black hole. We find $a^{*}$ to vary from 0.4 to 0.8. The value of the spin aligns with the earlier claim by Tchekhovskoy et al. (2014), where they predicted the spin of the source should be greater than $\geq 0.5$.
@@ IV.2 Coronal Variations
Figure 6: Optxagnf fitted spectral parameters variation with each other of Swift J1644+57 are presented. Panel (a) shows the variation of $N_{H}$ with respect to the normalized luminosity. The variation of $N_{H}$ with respect to $\Gamma$ is shown in panel (b). Panel (c) demonstrates the weak correlation between $\Gamma$ and scattering fraction $f_{pl}$. Variation of the $\Gamma$ and $\tau$ can be observed in panel (d). Panel (e) presents the variation of $R_{cor}$ and spin parameter. Uncorrelated variation between spin and $R_{cor}$ can be observed in panel (e).
To measure the $R_{cor}$ variation of Swift J1644+57, we employ the Optxagnf model to determine the spectral fits using the stacked Swift /XRT data. Through our spectral fitting, we find that at the peak luminosity of Swift J1644+57, the radial size of the corona was $31\pm 2.3~r_{\rm g}$. Subsequently, an almost similar size of the $R_{cor}$ was found in the XRT2 and XRT3 observational epochs. The size of the corona increases $35\pm 3.6~r_{\rm g}$ in the XRT4 epoch. Later on, in the XRT5 and XRT6 epochs, the $R_{cor}$ keeps increasing to $43\pm 3.7~r_{\rm g}$ and $51\pm 4.3~r_{\rm g}$, respectively. During the XRT7 observation, the $R_{cor}$ is $58\pm 4.1~r_{\rm g}$. Up to the XRT7 observation, we observed a sharp rise in the $R_{cor}$ . The fitted value of the $R_{cor}$ is $61\pm 4.2~r_{\rm g}$ at XRT8 and drops to $58\pm 4.5~r_{\rm g}$ at XRT10. From the XRT10 to XRT17 observations, the $R_{cor}$ varies little from $58\,r_{\rm g}$ to $64\,r_{\rm g}$, indicating a saturation level of the parameter. All estimated values of the coronal parameters are depicted in Fig. 8. In addition, we find through our analysis that the $R_{cor}$ is anti-correlated with the luminosity $\log(L/L_{\rm Edd})$ with the Pearson correlation coefficient: 0.87-0.87.
@@
$l_{I}$, $l_{C}$, and $l(t)$ are the initial, final, and the time-dependent Eddington ratios, respectively. Starting with an initial $R_{cor}$ $R_{I}$, the corona expands following the decay pattern of the luminosity profile and finally saturates to a value of $R_{C}$. We adopt $R_{I}=30\,r_{\rm g}$ and $R_{C}=65\,r_{\rm g}$. The normalized disk luminosity, $\log(L/L_{\rm Edd})$, at each observational epoch is displayed in Table 4. The predicted $R_{cor}$ at each epoch using Equation (8) is denoted with red circles in Fig. 8. The error in the theoretical values is a contribution arising from the observational error in the measurement of luminosity as presented in Table 4. We find that the evolution of theoretically speculated corona size agrees with that of the observational $R_{cor}$ parameter in Swift J1644+57, and considering the error bars, their variation remains similar.
@@
Figure 7: Markov Chain Monte-Carlo plot for parameter variations of XRT2 and XRT15 are presented. Goodman-Weare (EMCEE) (Foreman-Mackey et al., 2013) is opted for 32 walkers and a chain length of 1 million iterations, where the transient phase is removed by burning the chain up to 9984 iterations. The spin parameter was kept frozen at $a^{*}=0.60$ to perform the MCMCs.
@@
Figure 9: Edge-on view schematic diagram of $R_{cor}$ variation in the early and late stages of Swift J1644+57. The system comprises a disk around the SMBH, a corona on the disk, and a jet. The disk size is twice the tidal disruption radii, which is approximately estimated to be $2R_{\rm t}\sim 100\,r_{\rm g}$ for the $10^{6}M_{\odot}$ SMBH and a solar-type star (see equation 1).
@@ IV.2.1 Viability of TDE Corona Models
It should be noted that, to date, there exists no publicly available model in XSPEC to fit the hyper accretion regime of Swift J1644+57. The source was one of the very few jetted TDEs ever observed with a QPO feature. The initial accretion rate was $\sim 1000~L_{Edd}$ and was widely explored for its unique attributes. While choosing the model, one must be extremely cautious. Optxagnf is the only model in XSPEC that applies to the super-Eddington accretion regime of the AGNs, and measures the coronal radius parameter.
@@
Recently, Chakraborty et al. (2026) has proposed a coronal emission for TDEs (see their Fig. 7) where the optical depth could be higher, accompanied by a lower electron temperature. The parameters obtained from the Optxagnf model, presented in Table 4, exhibit a similar nature of corona. Although the morphology of the TDE corona could depart from the structure of the warm corona, the thermal properties, such as high $\tau$ and $kT_{s}$, resemble the warm corona of the Optxagnf model. The variation of $R_{cor}$ could be observed from the MCMC plots presented in the Fig. 7. However, it should be noted that Optxagnf, developed for AGNs, may not be physically self-consistent for hyper-Eddington jetted TDEs, such as for Swift J1644+57.
@@
Table 4: Optxagnf fitted spectral parameter are presented. We used stacked Swift/XRT over XMM-Newton as the former provides more degrees of freedom over the latter. The following parameters are kept frozen while fitting: $N_{H}^{Gal}=0.016\times 10^{22}\mathrm{cm}^{-2}$, model normalization $N_{optx}=1$, redshift $z=0.354$, distance $d=1402.8$ Mpc, and $log(r_{out})=5$.
ID MJD $N_{H}$ $\log(\frac{L}{L_{Edd}})$ $a^{*}$ $R_{cor}$ $kT_{s}$ $\tau$ $\Gamma$ $f_{pl}$ $dof/chi^{2}$
@@ V Summary
    By interpreting the hard X-ray-emitted region as a corona, we have estimated the size of the corona as a function of time. Considering the Optxagnf model, we confirm that the coronal radius parameter $R_{cor}$ increases over time for the observed period. We also find that the evolution of the corona size agrees with the theoretical conjecture made by Mummery and Balbus (2021).
@@
acknowledgments
The authors acknowledge an anonymous reviewer for their constructive comments. AC acknowledges the UPES Seed grant (UPES/R&D/SoAE/25062025/23) for partially supporting this research. The research of A.C., S.H., and S.S.H. were supported by the Canadian Space Agency (CSA) and the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grants and the Canada Research Chairs programs. The Basic Science Research Program has supported K.H. through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (2016R1A5A1013277 (K.H. and A.C.), 2020R1A2C1007219, and RS-2025-23323627 (K.H.)). K.H. thanks Andrew Mummery for fruitful discussions. AJ acknowledges support from the Fondecyt fellowship (Proyecto 3230303). NK acknowledge funds from European Union - Next Generation EU, Mission 4 Component 1 CUP C53D23001330006 and INAF Large Grant 2023 BLOSSOM F.O. 1.05.23.01.13. Research work at the Physical Research Laboratory, Ahmedabad, is funded by the Department of Space, Government of India. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France.